The position time graph of two cars A and B are straight lines making angles 30 degree and 60 degree with the timeaxis respectivily.The ratio of velocities of A and B is
1
Expert's answer
2019-07-03T08:52:20-0400

Velocity = Slope of the tangent at displacement-time diagram= tan(angle made of the tangent with time axis)

Velocity of car A = tan(30) = 1/1.732 = 0.577

Velocity of car B = tan(60)= 1.732

Ratio of Velocity A and Velocity B= 0.577/1.732 = 1/3= 0.33Velocity = Slope of the tangent at displacement-time diagram = tan(angle made of the tangent with time axis)

Velocity of car A = tan(30) = 1/1.732 = 0.577

Velocity of car B = tan(60)= 1.732

Ratio of Velocity A and Velocity B= 0.577/1.732 = 1/3= 0.33